a 81m high pole breaks but donot seperate and touch the ground 27m away from base at what height it will break from ground what is the solution through pythagoras theorem
Answer by stanbon(75887)

An 81m high pole breaks but does not separate, and touches the ground 27m away from the base. At what height does it break from ground? What is the solution using the Pythagorean equation.
--------------------------------------
Draw the picture.
You have a right triangle with base = 27
height = "h", and hypotenuse = "81-h"
------
Equation:
(81-h)^2 = h^2 + 27^2
-----
6561 - 162h + h^2 = h^2 + 729
----
-162h = -5832
---
height = 32 meters
==========================
Cheers,
Stan H.